If a plumber charges $35 for a house call plus $50 for every hour worked, write an equation that could be used to calculate the total cost of plumber's visit as a function of time.

Use your equation to find the total charges if the plumber worked for 2 1/2 hours?

Answer by kunal_122(19)

50h+35=cost
50*2.5+35=cost
125+35=cost
so cost =160